40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Sat Jan 7 00:45:47 2012, Frank, DailyProgress, NoiseBudget, current TTFSS noise budget TTFSS_enoise.pngTTFSS_fnoise.png
    Reply  Wed Jan 11 20:50:08 2012, tara, DailyProgress, NoiseBudget, current TTFSS noise budget 14_75MHz.png21_5MHz.png
       Reply  Tue Apr 23 18:43:39 2013, tara, DailyProgress, RefCav, HOM for new sideband frequencies HOM35.5MHz0.5m.pngHOM35.5MHz1m.pngHOM38MHz0.5m.pngHOM38MHz1m.pngshot_vs_beta.png
          Reply  Thu Jan 9 09:09:44 2014, Evan, DailyProgress, RefCav, Sideband frequencies for refcavs pdh_choice_close.pdfpdh_choice_close.pdf
Message ID: 781     Entry time: Sat Jan 7 00:45:47 2012     Reply to this: 782
Author: Frank 
Type: DailyProgress 
Category: NoiseBudget 
Subject: current TTFSS noise budget 

analyzed the data taken yesterday with the old TTFSS setup. Measured the EP noise (out1, common path, right after the mixer) and converted into frequency noise using an average number of the error signal slope of ~1MHz/V (which by the way is pretty bad).  Plot of electronic noise (demodulated PD signal) shows a noise level of ~20nV/rtHz, not super good but possible. With terminated input (~7nV/rtHz) it is slightly above the noise floor of the analyzer (6nV/rtHz)  which means it's about 4nV/rtHz which is OK.

TTFSS_enoise.png

If we compare the projected noise with the coating thermal noise we see that we are not able to measure anything useful with the current scheme. We need to gain about a factor of 10 to 20 more sensitivity. I suggest we switch to a resonant EOM to increase the modulation index and move to a lower frequency at the same time to increase the sensitivity of our RF PDs (old or new). This might also help us later when we need to stabilize the EOM temperature to minimize the RF-AM. HV-modulation might actually cause RF-AM which we are not able to cancel with our feedback from RF-AM sensing to EOM temp or DC actuator. We should check if we can use existing equipment like a 14.75MHz EOM we have or we can switch to the 21.5MHz which we used before. Switching can be done within one afternoon, two at most.

TTFSS_fnoise.png

 

We've moved on today towards a real TTFSS. Tara made the power cables and we quickly measured the noise of the EP of the new setup. Almost all harmonics are gone. Only tiny pickup of line harmonics is visible. Will make plots later.

ELOG V3.1.3-